OpenStudy (anonymous):

Find the absolute max/min values of : f(x)= (x-1)(x-7)^3+11 on the interval [1,4]. b.) interval [1,8] c.) [4,9]

OpenStudy (anonymous):

Is this a calculus problem?

OpenStudy (anonymous):

Yes it is

OpenStudy (anonymous):

Take derivative and second derivative. Where the derivative is zero will indicate a possible max or min. If the second derivative is positive, that value of x is a minimum; if negative, a maximum. Also, make sure to test the value of the function at the endpoints of the intervals, because the min or max might be there, even though the derivative isn't zero.

OpenStudy (anonymous):

Is that enough help?

OpenStudy (anonymous):

\[f(x)= (x-1)(x-7)^3+11\]\[f'(x)=(x-7)^3+3(x-1)(x-7)^2\]

OpenStudy (anonymous):

\[f"(x)=6(x-7)^2+6(x-1)(x-7)=12(x-7)(x-4)\]

OpenStudy (anonymous):

I see, I will try it from here. Thanks!

OpenStudy (anonymous):

My pleasure.